old family recipe makes 12 servings contains 2/3 cups of flour. if you have to make 27 servings how many cups of flour will you need.
Answer by algebrahouse.com(1659)

Set up a proportion:

12 servings / (2/3)cup = 27 servings / x cups
12x = (2/3)(27) {cross-multiplied}
12x = 18 {multiplied}
x = 18/12 {divided each side by 12}
x = 3/2 {reduced}

1 and 1/2 cups

For more help from me, visit: www.algebrahouse.com